OpenStudy (anonymous):

A. 2x + 8y = 10 B. 2x - 8y = 8 C. 2x + 8y = 16 D. 2x - 8y = 16

OpenStudy (anonymous):

Which is the equation of the line?

OpenStudy (anonymous):

standard equation y=mx+c the point where x=0 is called y intercept and represented by c so at x=0 y=2 implies c=2 y=mx+2 now at y=0 x=8 put it in equation 0=m*8+2 8m=-2 m=-2/8 m=-1/4

OpenStudy (anonymous):

equation y=-x/4+2 multiply by 8 on both side 8*y=-8*x/4+8*2 8y=-2x+16 2x+8y=16

OpenStudy (anonymous):

Thanks!